OpenStudy (anonymous):

in a large bag of candy,each of the 5 colors(red,orange,yellow,blue,and green) occurs with the same probability.you reach in and select 2 candies. find the probability that both candies are green A. 0.1 B. 0.2 C. O.4 D. 0.04

OpenStudy (anonymous):

@Deeunicorn3

OpenStudy (anonymous):

@shannon342

OpenStudy (anonymous):

idk wat to do now becuz now their both green

OpenStudy (mathmate):

What is the probability of picking a green candy at the first draw?

OpenStudy (anonymous):

is it 1 out of 5? or 1 out of 10?

OpenStudy (anonymous):

ok 2/5 is the probability again

OpenStudy (anonymous):

5 colors 2 possible outcomes of green

OpenStudy (anonymous):

Think of this a sequence of events... first you choose a green candy and then you choose a green candy again. Since there are a large number of candies in the bag, we'll assume that the probabilities don't change. So we're assuming independence. thus, P(G, G) = P(G)*P(G) = 1/5*1/5 = 1/25 = .04

OpenStudy (anonymous):

answer is D like @pgpilot326 explained

OpenStudy (anonymous):

ok thank u very much!